### v1

# Deblurring Neural Radiance Fields with Event-driven Bundle Adjustment

Abstract.
Neural Radiance Fields (NeRF) achieve impressive 3D representation learning and novel view synthesis results with high-quality multi-view images as input. However, motion blur in images often occurs in low-light and high-speed motion scenes, which significantly degrade the reconstruction quality of NeRF. Previous deblurring NeRF methods are struggling to estimate information during the exposure time, unable to accurately model the motion blur. In contrast, the bio-inspired event camera measuring intensity changes with high temporal resolution makes up this information deficiency. In this paper, we propose Event-driven Bundle Adjustment for Deblurring Neural Radiance Fields (EBAD-NeRF) to jointly optimize the learnable poses and NeRF parameters by leveraging the hybrid event-RGB data. An intensity-change-metric event loss and a photo-metric blur loss are introduced to strengthen the explicit modeling of camera motion blur. Experiment results on both synthetic data and real captured data demonstrate that EBAD-NeRF can obtain accurate camera poses during the exposure time and learn sharper 3D representations compared to prior works.
Neural radiance fields, event camera, image deblurring, novel view synthesis
ccs: Computing methodologies Shape representations
1. INTRODUCTION
Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) achieve 3D implicit representation learning and photo-realistic novel view synthesis results with high-quality 2D images and precise camera poses as input. In low-light scenes, a camera often requires a longer exposure time to obtain an image with sufficient brightness (Zamir et al., 2021b). A hand-held camera or a moving camera may cause motion blur in the captured image. Moreover, a high-speed moving camera can also cause motion blur, even in bright scenes with short exposure time. The blurred images will cause NeRF to learn a blurry 3D implicit representation, resulting in degraded quality of the synthesized novel view image. Thus, reconstructing sharp NeRF from blurry images is a practical problem for 3D reconstruction in low-light scenes (Yao et al., 2024).
Figure 1. Motivation of our proposed method. In a low-light scene or with a high-speed moving camera, motion blur usually occurs in the captured images. To reconstruct sharp NeRF from blurry images, Deblur-NeRF (Ma et al., 2022) uses a deformable sparse kernel to model the blur process. BAD-NeRF (Wang et al., 2023) linearly interpolates the camera poses and jointly learns the start and end poses of camera motion. However, these methods are unable to model complex motion blur and lack supervision information during exposure time with only a photo-metric loss. With the proposed event-driven bundle adjustment, our EBAD-NeRF can recover the accurate camera motion trajectory and consequently reconstruct sharper novel view images with sharper learned NeRF.
Motion blur is essentially caused by the camera pose change during exposure time. Recent deblurring NeRF works try to model the camera motion and optimize the estimated camera pose change with photo-metric loss (Ma et al., 2022; Wang et al., 2023; Qi et al., 2023). As shown in Figure 1, Deblur-NeRF (Ma et al., 2022) propose the deformable sparse kernel (DSK) module to simulate the change of rays from the camera. The camera motion trajectory of blur is implicitly learned in an additional multilayer perceptron (MLP), which is unstable when facing complex and severe motion blur. BAD-NeRF (Wang et al., 2023) uses linear interpolation to model the motion trajectory. The starting and ending poses of the trajectory are jointly optimized with the parameters of NeRF. This strategy works well for most motion blur, but a handheld camera is not necessarily shaking linearly under long exposures, and this could lead to a failure of reconstruction under the linear interpolation modeling. Although the previous works have tried different motion blur modeling methods for deblurring NeRF, relying solely on photo-metric loss cannot form effective constraints on the camera trajectory estimate. Because there is inherent information loss within the exposure time in the input blurred image, and the modeling is limited by this. The bio-inspired event camera can measure brightness changing asynchronously with high temporal resolution (Gallego et al., 2020), which can make up for the information loss of blurry images. Recently, E2NeRF (Qi et al., 2023) uses both event data and blurry images to model the blur in terms of image formation and achieved better results than the image-based deblurring works. However, camera poses of this work are obtained by pre-deblurred images and are not learnable, which limits its performance.
In this paper, we use learnable poses to model the camera motion trajectory. Event data is introduced into the camera pose learning by conducting an additional intensity-change-metric loss to effectively optimize the changes in adjacent poses, Unlike previous image-based works with only blurry images as supervision, which lacks constraints within the camera pose changes, we are able to fit the camera poses to the real motion trajectory during the blurring process and achieve sharper NeRF reconstruction. Specifically, for each view, we first use multiple learnable camera poses to model the camera motion trajectory and transfer them into the SE(3) space (Belta and Kumar, 2002; Wang et al., 2023). During the training, we jointly optimize these poses and the parameters of the NeRF network. A photo-metric blur loss is also used to supervise the modeling of blurry image formation. We introduce an intensity-change-metric event loss to supervise pose changes during the exposure time. The image formation is additionally strengthened by the event loss. The framework can estimate an accurate camera trajectory corresponding to the input blurry image, and a sharper NeRF is learned. We extend 5 blender synthetic scenes in Deblur-NeRF with event data and use a DAVIS-346 event camera (Taverni et al., 2018) to capture spatial-temporal aligned Event-RGB (ERGB) data with ground truth. Experiments on both synthetic data and real data demonstrate that our method can learn sharper neural radiation fields and more accurate camera motion trajectories compared with image deblurring methods and deblurring NeRF methods. To summarize, we present the following contributions:
1) A novel event-driven bundle adjustment deblurring neural radiance fields (EBAD-NeRF) framework is proposed to explicitly model image blur and jointly optimize the estimated camera motion trajectory and NeRF parameters.
2) An intensity-change-metric event loss and a photo-metric loss are presented to supervise the modeled motion blur in terms of both camera pose change and image formation conjugatively.
3) Experiments on both extended synthetic data and real captured data validate that our method achieves accurate motion trajectory estimation and high-quality 3D implicit reconstruction of the scene with severe blurry images and corresponding event data.
2. RELATED WORK
Table 1. Comparison of Previous Deblurring NeRF Works
Method RGB Event Motion Blur Model
Deblur-NeRF - 2D Blur Kernal
DP-NeRF - 3D Blur Kernal
Sharp-NeRF - Grid-Based Blur Kernal
BAD-NeRF - Linear Interpolation Bundle Adjustment
E2NeRF Poses from Pre-deblurred Images
EBAD-NeRF Event-driven Bundle Adjustment
2.1. Neural Radiance Fields
The neural radiation fields model achieves impressive novel view synthesis results and inspires a lot of subsequent research (Gao et al., 2022). For example, (Barron et al., 2021, 2022; Verbin et al., 2022; Xu et al., 2022) improve the quality of learned geometry and synthesized novel views. Other works try to improve the training and rendering speed such as FastNeRF (Garbin et al., 2021), MobileNeRF (Chen et al., 2023) and EfficientNeRF (Hu et al., 2022). PixelNeRF (Yu et al., 2021) and RegNeRF (Niemeyer et al., 2022) realize reconstructing high-quality NeRF with few input views. RawNeRF (Mildenhall et al., 2022) can recover a clean NeRF from noisy raw images. HDR-NeRF generates high dynamic novel view images with low dynamic images as input. Besides, nerf also requires accurate camera poses from COLMAP as input. BARF (Lin et al., 2021), L2G-NeRF, and Nope-NeRF (Bian et al., 2023) explore reconstructing NeRF when camera poses are not provided.
Blurry Images are often captured in low-light scenes or with a high-speed moving camera on drones or robots. Reconstructing sharp NeRF from blurry images also becomes a challenging problem. As shown in Table 1, Deblur-NeRF (Ma et al., 2022) proposes a deformable sparse kernel module to effectively model the blurring process. However, the blurring kernel is optimized based on 2D-pixel location independently. DP-NeRF (Lee et al., 2023) proposes a rigid blurring kernel and generates a 3D deformation field, which is constructed as the 3D rigid motion of the camera for each view. Sharp NeRF (Lee et al., 2024) proposed a learnable grid-based kernel to obtain sharp output from neural radiance fields. Inspired by BARF, BAD-NeRF (Wang et al., 2023) conducts linear interpolation for the camera poses during the exposure time and jointly optimizes the poses with bundle adjustment.
2.2. Image Deblurring
Traditional deblurring algorithms focus on finding the suitable blur kernel of a blurry image to recover a sharp image. Hand-crafted features or sparse priors are used to tackle this problem blur (Chan and Wong, 1998; Krishnan et al., 2011; Xu et al., 2013). Deep learning deblurring works directly learn an end-to-end mapping from blurry to sharp images (Tao et al., 2018a; Wieschollek et al., 2017; Zhang et al., 2020). MPRNet (Zamir et al., 2021a) and SRN (Tao et al., 2018b) achieve impressive single-image deblurring results. In addition to single-image deblurring, (Kim et al., 2016; Zhang and Carin, 2014; Zhang and Yang, 2015) proposes deblurring schemes for continuous video images, which used the characteristics of video images(such as temporal consistency) to better predict sharp images, but this also brought greater calculation amount. He et al. (He et al., 2016) weights a combination of input images to avoid explicit calculation of blur kernels and effectively improves processing speed. Other works have also tried implementing multiple image deblurring using convolutional neural network and Unet framework (Wieschollek et al., 2016; Su et al., 2017). Son et al.(Son et al., 2021) proposes an effective recurrent video deblurring network and achieves state-of-the-art performance. However, the information missing from blurry images during exposure time makes the robustness of the image-based method decreases in the face of various motion blur.
2.3. Event Camera
Event camera is based on a bio=inspired vision senor that measures brightness changing asynchronously (Gallego et al., 2020). The high temporal resolution and high dynamic visual data capturing paradigm makes event cameras ideal for flow estimation (Shiba et al., 2022; Hagenaars et al., 2021; Gehrig et al., 2021; Akolkar et al., 2020; Pan et al., 2020), feature detection and tracking(Xu et al., 2020; Zhang et al., 2021, 2022) and images deblurring (Pan et al., 2019; Lin et al., 2020; Shang et al., 2021; Xu et al., 2021). Pan et al. (Pan et al., 2019) propose a simple event-based double integral model (EDI) based on event generation to establish the connection between blurry images and events. Shang et al. (Shang et al., 2021), Jiang et al.(Jiang et al., 2020) and Sun et al. (Sun et al., 2022) using learning-based network to recover sharp images with events.
Recently, event-based NeRF works have emerged. Ev-NeRF (Hwang et al., 2023) models the measurement of the event sensor to learn grayscale neural radiance fields derived from the event stream. EventNeRF (Rudnev et al., 2023) reconstructs color NeRF with a color event camera. e-NeRF (Low and Lee, 2023) improves Ev-NeRF and EventNeRF for non-uniform camera motion. DE-NeRF (Ma et al., 2023) uses both RGB images and events to learn a deformable NeRF. E2NeRF first reconstructs a sharp NeRF with blurry images and corresponding event data and achieves state-of-the-art performance. However, it only explicitly models the blurring processing at the imaging aspect. For camera motion during the exposure time, E2NeRF relies on pre-deblurred images with EDI (Pan et al., 2019) algorithm and structure-from-motion method COLMAP for pose estimation. Notably, our proposed EBAD-NeRF transforms camera motion trajectories into a learnable model with event-driven bundle adjustment. The framework unlocks the potential of event information in explicit modeling of blur processing at the level of both image changes and camera motion.
Figure 2. Overview of our method. An image with motion blur is caused by the camera moving during the exposure time. We temporal evenly sample $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ and transfer them into SE(3) as learnable parameters. Then, the poses are input into the NeRF network to model the blurry image and event stream generation. The intensity-change-metric loss can optimize the modeling in terms of camera motion. The photo-metric loss can optimize and optimize the modeling in terms of image formation. With the joint learning of NeRF parameters $\theta$ and camera poses, a sharp NeRF is learned eventually.
3. METHOD
We introduce EBAD-NeRF to simultaneously learn the poses of the camera’s motion process and generate sharp neural radiance fields using both blurred images and the event stream within the corresponding exposure time. Events serve two key purposes within the framework: (1) optimizing camera trajectories within the exposure time, and (2) contributing to the physical formation of motion blur in the image by providing information on light intensity changes. Moreover, we employ a photo-metric blur loss to supervise blur formation in the RGB domain. Figure 2 shows the overview of our approach. Sec. 3.1 introduces the motion blur formation in the neural radiance fields. Sec. 3.2 illustrate the proposed event-driven bundle adjustment method. Sec. 3.3 shows the loss function and the derivation, and Sec. 3.4 describes the implementation details of our method.
3.1. Motion Blur Formation in Static 3D Scene
In a static 3D scene, camera motion blur is caused by the changes of camera pose $\mathbf{P}$ during the exposure time. A digital image $\mathcal{I}$ is obtained by measuring the integration of light intensity $I$ on the image sensor.
(1) $\mathcal{I}=\int_{t_{start}}^{t_{end}}I(t)\mathrm{d}t,\$
where ${t_{start}}$ and ${t_{end}}$ is the start and end time of exposure. The light intensity $I$ is only corresponding to the camera pose, and we can express $\mathbf{P}$ as a function of time $\mathbf{P}=p(t)$. Then, a blurry image $\mathcal{B}$ can be expressed as:
(2) $\mathcal{B}=\int_{t_{start}}^{t_{end}}I(P(t))\mathrm{d}t.\$
In NeRF (Mildenhall et al., 2021), the rendering results are closely related to camera poses. Given a camera pose $\mathbf{P}=P(t)$, for each pixel of the imaging plane, we can obtain a ray $\mathbf{r}$ that emits from the optical center of the camera and passes through the pixel $\mathbf{x}$. With stratified sampling, we can divide the part of this ray starting with $l_{near}$ and ending with $l_{far}$ into N equal parts and randomly sample points in each part. For each of the $N$ sampled points, we input the 3D coordination $\mathbf{o}$ in the world coordinate system and 2D view the direction $\mathbf{d}$ represented by the ray $\mathbf{r}$ into NeRF MLP $F_{\theta}$ with network parameters $\theta$:
(3) $(\mathbf{c},\sigma)=F_{\mathbf{\theta}}(\gamma_{o}(\mathbf{o}),\gamma_{d}(\mathbf{d})),$
the output of which is color $\mathbf{c}$ and density $\sigma$ and $\gamma(\cdot)$ is an encoder for a higher $K+1$ dimension of input of the network:
(4) $\gamma_{K}(x)=\{\sin(2^{k}\pi{x}),\cos(2^{k}\pi{x})\}_{k=0}^{K}.$
Then, we can obtain $N$ colors $\{\mathbf{c}_{i}\}_{i=1}^{N}$ and densities $\{\mathbf{\sigma}_{i}\}_{i=1}^{N}$ of the sampled points. By conducting volume rendering:
(5) $C(\mathbf{r},\mathbf{x})=\sum_{i=1}^{N}T_{i}(1-\exp(-{\sigma}_{i}{\delta}_{i}))\mathbf{c}_{i},$
$\text{where}\quad T(i)=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}),$
we can get the final color value of the pixel $\mathbf{x}$ passed by the ray $\mathbf{r}$ corresponding to the pose $\mathbf{P}$. $\delta_{i}=l_{i+1}-l_{i}$ is the distance between adjacent sampled points, and $T_{i}$ is the transparency between $l_{n}$ and sampled point.
If we assume that the learned NeRF MLP is sharp, we can model the image motion blur by discretizing Eq. (2) with $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ sampled evenly over the exposure time:
(6) ^$\hat{\mathcal{B}}(\mathbf{x})=\frac{1}{p}\sum_{i=1}^{p}C(P(t),\mathbf{x}),\mathbf{x}\in\mathcal{X},\$
where $\mathcal{X}$ represents the pixels of the image sensor. Since we use virtual sharp image color $C(P(t),\mathbf{x})$ to replace the light intensity $I(P(t))$, we need to multiply each item by a weight of time in Eq. (6). As the poses are temporally evenly sampled, we can directly use an average $\frac{1}{p}$ to represent the weight.
At this point, we have established the connection between image motion blur, camera pose, and neural radiation field.
3.2. Event-driven Bundle Adjustment
In BAD-NeRF (Wang et al., 2023), only the start pose $\mathbf{p}_{1}$ and end pose $\mathbf{p}_{p}$ of the exposure time is parameterized and learned, and the motion process is represented with a linear model. This approach may fail in long exposure or high-speed situations where the camera is likely to have complex motion (non-uniform motion, curved motion) during exposure time. In order to better fit the camera motion trajectory, we express the camera poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ as $\{\mathbf{T}_{i}\}_{i=1}^{p}\in\mathbf{SE}(3)$.
At this time, in addition to recovering the starting and ending poses, the motion trajectory modeling must also find the relationship between adjacent poses. We can model this correlation by modeling the generation of events, which measures the light intensity change in the log domain asynchronously:
(7) $E(t_{1},t_{2})=\left\{\begin{aligned} \lfloor\frac{\log(I(t_{2}))-\log(I(t_{1})))}{\Theta}\rfloor,&I(t_{2})\leq I(t_{1})\\ \lceil\frac{\log(I(t_{2}))-\log(I(t_{1}))}{\Theta}\rceil,&I(t_{2})>I(t_{1})\end{aligned}.\right.$
Note that we round the modeled events because an event is only triggered when the light intensity change threshold reaches $\Theta$.
According to previous event simulation work, V2E (Hu et al., 2021), we can convert the color of the images into grayscale and model the event generation. Based on this, the estimated event data can be expressed by bringing the predicted color in Eq. (5) into Eq. (7):
(8) $\hat{E}(t_{1},t_{2})=\left\{\begin{aligned} \lfloor\frac{\log(\overline{C}(T({t_{2}})))-\log(\overline{C}(T(t_{1})))}{\Theta}\rfloor,&\overline{C}(t_{2})\leq\overline{C}(t_{1})\\ \lceil\frac{\log(\overline{C}(T({t_{2}})))-\log(\overline{C}(T(t_{1})))}{\Theta}\rceil,&\overline{C}(t_{2})>\overline{C}(t_{1})\end{aligned},\right.$
where $\overline{C}(\cdot)$ means taking the average on the RGB channels to get the intensity value.
The computation of $\hat{E}(\mathbf{x},t_{1},t_{2})$ is differentiable with respect to $\{\mathbf{T}_{i}\}_{i=1}^{p}$, and we can jointly optimize the sampled camera poses and parameters of $F_{\theta}$ by using actual events as supervision.
3.3. Loss Functions
We introduce the intensity-change-metric event loss to optimize the camera motion poses $\{\mathbf{T}_{i}\}_{i=1}^{p}$ during exposure time as well as the NeRF parameter $\theta$ by minimizing the differences between the numbers of predicted and actual events on all pixels $\mathbf{x}\in\mathcal{X}$:
(9) $\mathcal{L}_{event}=\frac{1}{p}\sum_{k=1}^{p-1}\|(\hat{E}_{k}(t_{k},t_{k+1}))-(E_{k}(t_{k},t_{k+1}))\|_{2}^{2}.$
In the training phase, we only calculate the event loss between adjacent sampling poses because if the sampling frames are far apart, the event loss will fluctuate greatly, thus affecting the stability of the network.
We also conduct the photo-metric blur loss between the predicted and the input blurry images:
(10) $\mathcal{L}_{blur}={\|\hat{B}(\mathbf{x})-B(\mathbf{x})\|}_{2}^{2},$
which also has a constraint on both the camera poses and neural radiance fields, as proved in BAD-NeRF. To strengthen the learning, the final loss is the sum of $\mathcal{L}_{event}$ and $\mathcal{L}_{blur}$:
(11) $\mathcal{L}_{blur}=\lambda\mathcal{L}_{event}^{c}+\mathcal{L}_{blur}^{c}+\mathcal{L}_{blur}^{f},$
where $\lambda$ is a weight parameter of event loss. We used the design of the fine and coarse network in NeRF to conduct blur loss $\mathcal{L}^{f}_{blur}$ and $\mathcal{L}^{c}_{blur}$ on both of the networks. Event loss $\mathcal{L}^{f}_{event}$ is only calculated for the fine network because it can render enough texture details for event estimation.
The learnable camera poses $\{\mathbf{T}_{i}\}_{i=1}^{p}$ and NeRF parameter $\theta$ are optimized by the Jacobians of event loss:
(12) $\frac{\partial{\mathcal{L}_{event}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\mathcal{E}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{E}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
(13) $\frac{\partial{\mathcal{L}_{event}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{event}}{\partial\mathcal{E}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{E}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}},$
and the Jacobians of blur loss:
(14) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\mathcal{B}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{B}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
(15) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{blur}}{\partial\mathcal{B}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{B}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}},$
where $m$ is the number of pixels of the image.
Then, the network can learn an accurate camera motion and sharp 3D implicit representation with gradient propagation.
3.4. Implementation Details
We train the EBAD-NeRF on a single NVIDIA RTX 3090 GPU. The training time has not increased compared with BAD-NeRF because there is no extra rendering computation for event loss. The sample number of the rays $N=64$ for fine and coarse networks. The number of sampled poses of each view $p$ is set to 4, and the weight of event loss $\lambda$ is set to 0.005. Sec.4.4 evaluates the influence of the parameters. Additionally, we use $\Theta=0.3$ as the threshold of the event generation, which is a typical value in event-based vision (Gallego et al., 2020). A coarse initial pose for each view is given at the start of training.
4. EXPERIEMENTS
4.1. Datasets
4.1.1. Synthetic Data:
We extend five Blender scenarios of Deblur-NeRF (Cozyroom, Factory, Pool, Tanabata, and Wine). (1) We increase the camera shake amplitude of the Factory, Pool, Tanabata, and Wine scenes to increase the blur degree of the training images and maintain the camera shaking in Deblur-NeRF of the Cozyroom scene to evaluate the effect of different shaking scales. (2) During the camera motion, we added the speed changes often occurring when the hand-held camera shakes. (3) To generate corresponding event data, we input the virtual sharp frames of the camera shake process rendered by Blender into the event simulation tool V2E (Hu et al., 2021) to generate the corresponding event stream.
4.1.2. Real Data:
To test the effectiveness of our method on real captured data, we captured two sets of real data using the DAVIS-346 (Taverni et al., 2018) event camera, which can capture spatial-temporal aligned event data and RGB data at the same time. Both sets of data are captured in low-light scenes, so a longer exposure time (100ms) is needed for RGB data to get a bright enough image. We use a handheld camera to capture training data with blur and a tripod to mount the camera to capture sharp testing data. Each scene contains 16 views of blurry images and corresponding events for training and 4 novel view sharp images for training.
4.2. Comparing Method
4.2.1. Image Deblurring Method:
For comparison, we selected two classic learning-based image deblurring methods, MPR (Zamir et al., 2021a) and SRN (Tao et al., 2018b). In order to make a more fair comparison in terms of data modalities, we selected the D2Net method, which also uses event and image modal data fusion to achieve image deblurring. We input the images deblurred by these three methods into the original NeRF for the novel view generation task to learn a sharp 3D representation.
4.2.2. Deblurring NeRF Method:
We selected image-based deblurring NeRF method Deblur-NeRF (Ma et al., 2022) and BAD-NeRF (Wang et al., 2023) and ERGB-dual-modal-based deblurring NeRF method E2NeRF (Qi et al., 2023) for comparison. For Deblur-NeRF, we use the default parameters for the spare deformation kernel. For BAD-NeRF, we use the number of interpolated virtual cameras as 7 to get sufficiently satisfactory results, as described in the paper. For E2NeRF, we follow the paper and use $b=4$ to train the model. All NeRF-based methods are trained with 5000 rays of batch size and 64 sampling points for coarse and fine networks. We perform 100,000 iterations on synthetic data. On real data, due to the reduction in image resolution and the number of views, 50,000 iterations are enough for all NeRF-based methods to converge to good enough results.
4.2.3. Metrics:
We use PSNR, SSIM, and LPIPS (Zhang et al., 2018) to evaluate the NeRF reconstruction results and use absolute trajectory error (ATE) to evaluate the fitting accuracy of the camera motion trajectory (Grupp, 2017).
Table 2. Quantitative Ablation Study on the Tanabata Scene.
PSNR↑ SSIM↑ LPIPS↓ ATE
BAD-NeRF 20.65 .7567 .3349 0.0514 $\pm$ 0.0204
EBAD-NeRF-noe 22.00 .7961 .3728 0.0462 $\pm$ 0.0183
EBAD-NeRF-linear 23.75 .8478 .2881 0.0418 $\pm$ 0.0160
EBAD-NeRF-full 24.99 .8800 .2069 0.0301 $\pm$ 0.0122
Figure 3. Qualitative Ablation Study on the Tanabata Scene. EBAD-NeRF-linear and EBAD-NeRF-noe are defined in Sec. 4.3. With event-driven bundle adjustment, the rendering results of learned NeRF are sharper and clearer.
Figure 4. Ablation study of trajectory fitting accuracy on Tanabata scene. EBAD-NeRF-linear and EBAD-NeRF-noe are defined in Sec. 4.3. The figure shows that the event-driven bundle adjustment significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than other methods.
Table 3. Quantitative Results on Deblurring Views of Blender Scenes. The Best Results are Shown in Bold.
Cozyroom Factory Pool Tanabata Wine Average
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
NeRF 27.04 .9034 .2875 20.54 .6776 .5233 27.39 .8352 .4356 19.13 .6691 .5743 20.11 .6732 .5832 22.84 .7517 .4808
D2Net 28.25 .9271 .2083 21.15 .7096 .4428 28.23 .8589 .3537 19.34 .6947 .4973 20.60 .6977 .4894 23.52 .7776 .3983
NeRF+D2Net 28.11 .9234 .2256 21.20 .7100 .4485 28.23 .8576 .3757 19.48 .6964 .5077 20.64 .6976 .5097 23.53 .7770 .4134
MPR 26.08 .9025 .2505 21.07 .6985 .4503 27.09 .8318 .3692 19.42 .7046 .4983 20.31 .6907 .5172 22.80 .7656 .4171
NeRF+MPR 26.78 .9094 .2642 21.14 .7023 .4660 27.33 .8388 .4063 19.79 .7131 .5161 20.56 .6936 .5504 23.12 .7714 .4406
SRN 28.11 .9216 .1943 22.96 .7752 .3384 28.58 .8705 .2882 19.80 .7202 .4208 21.30 .7297 .4160 24.15 .8034 .3315
NeRF+SRN 28.29 .9271 .2073 23.22 .7900 .3309 28.88 .8812 .2964 20.11 .7340 .4238 21.61 .7448 .4215 24.42 .8154 .3360
Deblur-NeRF 28.49 .9275 .1993 21.67 .7213 .4431 27.99 .8581 .3527 18.99 .6685 .4843 20.65 .6915 .4914 23.56 .7734 .3942
BAD-NeRF 28.80 .9278 .1872 20.39 .6696 .4052 29.46 .8867 .2416 20.35 .7475 .3363 22.10 .7540 .3831 24.22 .7971 .3107
E2NeRF 30.17 .9459 .1057 27.90 .9046 .2638 29.29 .8841 .2420 24.02 .8625 .2906 25.63 .8688 .2919 27.40 .8932 .2388
EBAD-NeRF 30.53 .9475 .1120 28.10 .9085 0.159 31.50 .9193 .1607 24.91 .8783 .2071 26.66 .8732 .2230 28.30 .9031 .1749
4.3. Event-driven Bundle Adjustment Analysis
We analyze the proposed event-driven bundle adjustment at two levels: 3D reconstruction quality and motion trajectory fitting accuracy. As shown in Table 2, Figure 3 and Figure 4, EBAD-NeRF-noe training proposed EBAD-NeRF without event data as supervision. EBAD-NeRF-linear represents the results of adding event data constraints to BAD-NeRF directly.
4.3.1. Defect of Linear Interpolation Bundle Adjustment:
The quantitative results in Table 2 demonstrate that without event enhancement, the bundle adjustment constrained by linear interpolation (BAD-NeRF) is even worse than the unconstrained bundle adjustment (EBAD-NeRF-noe) for fitting irregular camera motions. The same conclusion will be obtained by comparing EBAD-NeRF-linear and EBAD-NeRF-full, both with event enhancement.
4.3.2. Effect of Linear Event in Bundle Adjustment:
Comparing BAD-NeRF and EBAD-NeRF-linear, even with the negative impact of linear interpolation bundle adjustment in Table 2, EBAD-NeRF-linear still has improvements in trajectory fitting and reconstruction results. Comparing EBAD-NeRF-linear and EBAD-NeRF under the constraints of event information, event-driven bundle adjustment achieved the best trajectory fitting and reconstruction results.
To sum up, during the NeRF training, event data can not only provide image-level optimization information but also pose-level practical constraints. Our proposed EBAD-NeRF effectively utilizes the ERGB-dual-modal data to achieve this goal. The qualitative comparison in Figure 3 and Figure 4 also testify to this.
Figure 5. The influence of the number of learnable poses $p$ and the weight parameter $\lambda$ on the Cozyroom scene. The results are average for reconstructed images from both deblurring and novel views.
4.4. Analysis on Parameters
We evaluate the rendering results on the Cozyroom scene to analyze the superparameters $p$ and $\lambda$ of the proposed network. The red, green, and blue lines in Figure 5 represent PSNR, SSIM, and LPIPS, respectively.
4.4.1. The influence of $b$:
Intuitively, the more virtual sharp frames, the more accurate the simulation of camera motion blur, and the corresponding network learning results will be better. The results in Figure 5 are consistent with this. When the number of virtual frames gradually increases from 1, the performance of the network also becomes significantly better. After exceeding 5, the number of virtual frames has almost no impact on the reconstruction results. Therefore, $p=5$ is set in our experiments.
4.4.2. The influence of $\lambda$:
As shown in Figure 5, $\lambda$ has the best effect when set to 0.005, and if the weight of event loss is further increased, the network performance will gradually decrease.
Table 4. Quantitative Results on Novel Views of Blender Scenes. The Best Results are Shown in Bold
Cozyroom Factory Pool Tanabata Wine Average
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
NeRF 26.98 .9021 .2875 20.58 .6953 .5226 27.22 .8345 .4320 19.12 .6882 .5654 19.91 .6763 .5852 22.76 .7593 .4785
NeRF+D2Net 28.02 .9219 .2264 21.49 .7345 .4490 28.04 .8582 .3745 19.70 .7181 .4990 20.43 .7137 .5092 23.54 .7893 .4116
NeRF+MPR 26.60 .9066 .2665 22.08 .7355 .4629 27.20 .8368 .4045 19.81 .7183 .5112 20.45 .6968 .5513 23.23 .7788 .4393
NeRF+SRN 28.21 .9251 .2076 23.18 .8018 .3300 28.78 .8808 .2939 21.03 .7708 .4133 21.99 .7637 .4203 24.64 .8284 .3330
Deblur-NeRF 28.42 .9260 .2000 22.12 .7490 .4369 27.92 .8594 .3491 20.77 .7412 .4605 21.13 .7128 .4859 24.07 .7977 .3865
BAD-NeRF 28.52 .9250 .1918 21.11 .7169 .4077 29.70 .8919 .2376 22.37 .8102 .3267 22.79 .7791 .3812 24.90 .8246 .3090
E2NeRF 30.07 .9469 .1062 27.78 .9090 .2654 29.34 .8905 .2374 24.25 .8730 .2882 25.70 .8772 .2910 27.43 .8993 .2376
EBAD-NeRF 30.60 .9476 .1138 28.25 .9146 .1583 31.62 .9244 .1559 25.45 .8898 .2056 26.72 .8889 .1961 27.80 .9131 .1784
Table 5. Quantitative Results on Real Scenes. The Results are the Average of the Scenes, and The Best Results are Shown in Bold.
NeRF NeRF+D2Net NeRF+MPR NeRF+SRN Deblur-NeRF BAD-NeRF E2NeRF EBAD-NeRF
PSNR↑ 24.87 26.31 24.35 25.91 25.59 28.10 27.78 29.60
SSIM↑ .8562 .8829 .8539 .8706 .8744 .8839 .9062 .9175
LPIPS↓ .5005 .4092 .4963 .4198 .3885 .3632 .2403 .2156
4.5. Quantitative Results
4.5.1. Quantitative Results of Reconstruction on Blender Scenes:
In Tables 3 and 4, we evaluate the reconstruction results of learned neural radiance fields in five blender scenarios from deblurring and novel views, respectively. SRN and SRN+NeRF achieve the best results in single-image-based deblurring methods. With the help of event data, D2Net is also slightly better than MPR. For Deblur-NeRF and BAD-NeRF, although they have been further improved compared to NeRF+D2Net, NeRF+MPR, and NeRF+SRN, they still cannot surpass event-enhanced E2NeRF and our proposed EBAD-NeRF. Note that there is no apparent difference between the E2NeRF and EBAD-NeRF on the Cozyroom scene where linear camera shake produces motion blur, and E2NeRF is even better than EBAD-NeRF on LPIPS. However, in the other four scenes with more complex camera shaking as described in Sec. 4.1, the results of EBAD-NeRF in both deblurring views and novel views are significantly better than E2NeRF, which indicates that lacking the optimization of poses limits the role of events in 3D reconstruction in E2NeRF, especially when facing complex camera motion. Accordingly, EBAD-NeRF uses event-driven bundle adjustment to jointly optimize the motion poses of the camera, further releasing the potential of event data in reconstructing a sharp NeRF with blurry images.
Table 6. Absolute Trajectory Error Results of Blender Scenes.
ATE
COLMAP-Blur (NeRF (Mildenhall et al., 2021)) 0.0476 $\pm$ 0.0171
COLMAP-EDI (E2NeRF (Qi et al., 2023)) 0.0419 $\pm$ 0.0159
Linear Bundle Adjustment (BAD-NeRF (Wang et al., 2023)) 0.0462 $\pm$ 0.0142
Event-driven Bundle Adjustment (Ours) 0.0383 $\pm$ 0.0143
4.5.2. Quantitative Results of Pose Estimation on Blender Scenes:
In Table 6, we use absolute trajectory error to evaluate the accuracy of camera motion trajectory estimation by different methods. The results in the table are the average of five blender scenes. COLMAP-Blur means directly inputting the blurry images into COLMAP to estimate the camera poses of each view. This is a typical pose estimation method in NeRF-related work. COLMAP-EDI represents the method used in E2NeRF. It uses the EDI (Pan et al., 2019) algorithm to first use events to deblur the image and then input the sharp image sequence into COLMAP to estimate camera poses of the blur process. Linear bundle adjustment represents the linear interpolation method used by BAD-NeRF to jointly learn the pose changes of each view. Our proposed event-driven bundle adjustment method introduces event into joint learning of camera pose changes and sharp NeRF, achieving the best results in restoring the camera pose, which also promotes EBAD-NeRF to further improve the reconstruction quantification results as in Tables 3 and 4.
4.5.3. Quantitative Reconstruction Results on Blender Scenes:
To verify the effectiveness of our method on real data, we conducted experiments with two real scenes captured by DAVIS-346. As shown in Table 5, the comparison results with other methods are highly consistent with the results on blender scenes in Tables 3 and 4, which verifies the reliability of our method in practical applications.
Figure 6. Qualitative results of deblurring views on blender Pool scene.
Figure 7. Qualitative results of novel views on blender Wine and Factory scenes (first and second row), real scene (third row).
4.6. Qualitative Results
4.6.1. Qualitative Results of Reconstruction on Blender Scenes
As shown in Figure 6 and the first and second row in Figure 7, we compared the results of the synthesis deblurring view on the pool scene and novel view on the Wine and Factory scenes, respectively. For the deblurring view results, the results of EBAD-NeRF are sharper than other state-of-the-art single-image deblurring methods, event-RGB-based deblurring methods, and deblurring NeRF methods. For the green plants and wooden boxes in the scene, some artifacts appear in the results of E2NeRF, and our method effectively recovers the texture details. In the Wine and Factory scenes, though E2NeRF realizes noticeable results for the characters and letters, our method is closer to ground truth in color, clarity, and line thickness.
Figure 8. Qualitative evaluation of trajectory fitting accuracy on Tanabata scene. The figure shows that the event-driven bundle adjustment significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than other methods. Note that though there is event enhancement on blurry image formation in E2NeRF, the motion trajectory is distorted and leads to the inaccurate NeRF.
4.6.2. Qualitative Results of Pose Estimation on Blender Scenes
In Figure 8, we compare the estimated camera trajectory of different methods from the three dimensions of roll, pitch, and yaw. The purple curve representing EBAD-NeRF is the closest to the dashed line representing ground truth, which is consistent with the quantitative data in Table 6.
4.6.3. Qualitative Results on Real Scenes
As in the third row of Figure 7, our method better recovers the white letters on the blackboard, compared to the pose-fixed event-RGB-based method E2NeRF, while BAD-NeRF is limited in reconstructing smooth areas (white areas on the podium). The results of other image-deblurring-NeRF methods and Deblur-NeRF are all limited by the severe blurry input, which is the same as the quantitative results in Table 5.
Experiments on synthetic and real data show that our method can learn better poses of the camera motion blur process with event-driven bundle adjustment compared to the previous deblurring NeRF method. On this basis, the event data is also superimposed on the image blur process supervision, eventually achieving better 3D reconstruction and novel view synthesis effects.
5. Conclusion
In conclusion, this paper presents a novel approach, leveraging event-driven bundle adjustment, to address the challenge of modeling camera motion blur within neural radiance fields (NeRFs). Our framework draws inspiration from the emerging field of event-based vision, which offers high temporal resolution and asynchronous events, ideal for capturing dynamic scenes with motion blur. By integrating intensity-change-metric event loss and photo-metric blur loss, our framework enables the simultaneous optimization of blur modeling alongside NeRF reconstruction. Experiments on synthetic and real-world datasets demonstrate the efficacy of our approach in accurately estimating camera poses and producing sharp NeRF reconstructions.

### v2

@@
 Abstract.
-Neural Radiance Fields (NeRF) achieve impressive 3D representation learning and novel view synthesis results with high-quality multi-view images as input.
-However, motion blur in images often occurs in low-light and high-speed motion scenes, which significantly degrade the reconstruction quality of NeRF.
-Previous deblurring NeRF methods are struggling to estimate information during the exposure time, unable to accurately model the motion blur.
-In contrast, the bio-inspired event camera measuring intensity changes with high temporal resolution makes up this information deficiency.
+Neural Radiance Fields (NeRF) achieves impressive 3D representation learning and novel view synthesis results with high-quality multi-view images as input.
+However, motion blur in images often occurs in low-light and high-speed motion scenes, which significantly degrades the reconstruction quality of NeRF.
+Previous deblurring NeRF methods struggle to estimate pose and lighting changes during the exposure time, making them unable to accurately model the motion blur.
+The bio-inspired event camera measuring intensity changes with high temporal resolution makes up this information deficiency.
 In this paper, we propose Event-driven Bundle Adjustment for Deblurring Neural Radiance Fields (EBAD-NeRF) to jointly optimize the learnable poses and NeRF parameters by leveraging the hybrid event-RGB data.
 An intensity-change-metric event loss and a photo-metric blur loss are introduced to strengthen the explicit modeling of camera motion blur.
-Experiment results on both synthetic data and real captured data demonstrate that EBAD-NeRF can obtain accurate camera poses during the exposure time and learn sharper 3D representations compared to prior works.
+Experiments on both synthetic and real-captured data demonstrate that EBAD-NeRF can obtain accurate camera trajectory during the exposure time and learn a sharper 3D representations compared to prior works.
 Neural radiance fields, Event camera, Image deblurring, Novel view synthesis
-ccs: Computing methodologies Shape representations
+copyright: noneccs: Computing methodologies Computer vision
 1. Introduction
-Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) achieve 3D implicit representation learning and photo-realistic novel view synthesis results with high-quality 2D images and precise camera poses as input.
-In low-light scenes, a camera often requires a longer exposure time to obtain an image with sufficient brightness (Zamir et al., 2021b).
-A hand-held camera or a moving camera may cause motion blur in the captured image.
+Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) achieves 3D implicit representation learning and photo-realistic novel view synthesis results with high-quality 2D images and precise camera poses as input.
+In low-light scenes, a camera often requires a longer exposure time to obtain an image with sufficient brightness (Zamir et al., 2021b), and a handheld camera may cause motion blur in the captured image.
 Moreover, a high-speed moving camera can also cause motion blur, even in bright scenes with short exposure time.
-The blurred images will cause NeRF to learn a blurry 3D implicit representation, resulting in degraded quality of the synthesized novel view image.
-Thus, reconstructing sharp NeRF from blurry images is a practical problem for 3D reconstruction in low-light scenes (Yao et al., 2024).
+The blurred images will cause NeRF to learn a blurry 3D implicit representation, resulting in degraded quality of the synthesized novel view images.
+Thus, it is a practical problem to reconstruct a sharp NeRF from blurry images when facing low-light and high-speed scenes (Yao et al., 2024).
 Figure 1.
@@
 However, these methods are unable to model complex motion blur and lack supervision information during exposure time with only a photo-metric loss.
-With the proposed event-driven bundle adjustment, our EBAD-NeRF can recover the accurate camera motion trajectory and consequently reconstruct sharper novel view images with sharper learned NeRF.
-Motion blur is essentially caused by the camera pose change during exposure time.
-Recent deblurring NeRF works try to model the camera motion and optimize the estimated camera pose change with photo-metric loss (Ma et al., 2022; Wang et al., 2023; Qi et al., 2023).
-As shown in Figure 1, Deblur-NeRF (Ma et al., 2022) propose the deformable sparse kernel (DSK) module to simulate the change of rays from the camera.
-The camera motion trajectory of blur is implicitly learned in an additional multilayer perceptron (MLP), which is unstable when facing complex and severe motion blur.
-BAD-NeRF (Wang et al., 2023) uses linear interpolation to model the motion trajectory.
-The starting and ending poses of the trajectory are jointly optimized with the parameters of NeRF.
-This strategy works well for most motion blur, but a handheld camera is not necessarily shaking linearly under long exposures, and this could lead to a failure of reconstruction under the linear interpolation modeling.
-Although the previous works have tried different motion blur modeling methods for deblurring NeRF, relying solely on photo-metric loss cannot form effective constraints on the camera trajectory estimate.
-Because there is inherent information loss within the exposure time in the input blurred image, and the modeling is limited by this.
-The bio-inspired event camera can measure brightness changing asynchronously with high temporal resolution (Gallego et al., 2020), which can make up for the information loss of blurry images.
-Recently, E2NeRF (Qi et al., 2023) uses both event data and blurry images to model the blur in terms of image formation and achieved better results than the image-based deblurring works.
-However, camera poses of this work are obtained by pre-deblurred images and are not learnable, which limits its performance.
-In this paper, we use learnable poses to model the camera motion trajectory.
-Event data is introduced into the camera pose learning by conducting an additional intensity-change-metric loss to effectively optimize the changes in adjacent poses, Unlike previous image-based works with only blurry images as supervision, which lacks constraints within the camera pose changes, we are able to fit the camera poses to the real motion trajectory during the blurring process and achieve sharper NeRF reconstruction.
+With the proposed event-driven bundle adjustment, our EBAD-NeRF can leverage the event to recover the accurate camera trajectory and learn a sharper NeRF, resulting in sharper novel view image rendering than the previous methods.
+For static scenes, motion blur is caused by the camera pose change during the exposure time.
+The core problem of reconstructing a sharp NeRF is accurately restoring the motion trajectory to better model the motion blur formation.
+Correspondingly, we need to establish a connection between scene radiance fields and camera pose change.
+Recent deblurring NeRF works try to resolve this by using a learnable blur kernel (Ma et al., 2022; Lee et al., 2023, 2024) or interpolating the camera poses with learnable poses at the start and end of the exposure (Wang et al., 2023).
+However, they solely implicitly model the motion and are supervised by RGB images, which is insufficient to establish a strong connection between camera pose change and NeRF parameters.
+This results in unstable network training and a decrease in performance for complex and severe motion blur.
+As in Figure 1, Deblur-NeRF (Ma et al., 2022) and Bad-NeRF (Wang et al., 2023) all learn an inaccurate motion trajectory and an inferior NeRF with blurry rendering results.
+The bio-inspired event camera can measure brightness change asynchronously with high temporal resolution (Gallego et al., 2020), which makes up for the information loss in blurry images.
+Recently, E2NeRF (Qi et al., 2023) uses the event data and blurry images to enhance the model of motion blur in terms of image formation and achieves better results than the image-based deblurring works.
+However, it uses pre-deblurred images and COLMAP (Schonberger and Frahm, 2016) to estimate camera poses, which is not learnable during training, and the pose change information in the event data is not fully modeled and utilized.
+In this paper, we model the motion blur in terms of both motion trajectory and camera imaging.
+Event data is introduced to establish an explicit connection between the actual camera motion trajectory and the target sharp NeRF by conducting an intensity-change-metric event loss.
 Specifically, for each view, we first use multiple learnable camera poses to model the camera motion trajectory and transfer them into the SE(3) space (Belta and Kumar, 2002; Wang et al., 2023).
 During the training, we jointly optimize these poses and the parameters of the NeRF network.
+We introduce an intensity-change-metric event loss to supervise pose changes during the exposure time.
 A photo-metric blur loss is also used to supervise the modeling of blurry image formation.
-We introduce an intensity-change-metric event loss to supervise pose changes during the exposure time.
-The image formation is additionally strengthened by the event loss.
-The framework can estimate an accurate camera trajectory corresponding to the input blurry image, and a sharper NeRF is learned.
-We extend 5 blender synthetic scenes in Deblur-NeRF with event data and use a DAVIS-346 event camera (Taverni et al., 2018) to capture spatial-temporal aligned Event-RGB (ERGB) data with ground truth.
-Experiments on both synthetic data and real data demonstrate that our method can learn sharper neural radiation fields and more accurate camera motion trajectories compared with image deblurring methods and deblurring NeRF methods.
+The image formation is also additionally strengthened by the event loss.
+We extend 5 blender synthetic scenes in Deblur-NeRF with event data and use a DAVIS-346 event camera (Taverni et al., 2018) to capture spatial-temporal aligned Event-RGB (ERGB) real data with ground truth.
+Experiments on both synthetic and real data demonstrate that our method can learn superior NeRF and more accurate camera motion trajectories than previous image-based or ERGB-based deblurring NeRF methods.
+The rendering results are also better than the state-of-the-art image-based or ERGB-based image deblurring methods.
 To summarize, we present the following contributions:
 1) A novel event-driven bundle adjustment deblurring neural radiance fields (EBAD-NeRF) framework is proposed to explicitly model image blur and jointly optimize the estimated camera motion trajectory and NeRF parameters.
-2) An intensity-change-metric event loss and a photo-metric loss are presented to supervise the modeled motion blur in terms of both camera pose change and image formation conjugatively.
-3) Experiments on both extended synthetic data and real captured data validate that our method achieves accurate motion trajectory estimation and high-quality 3D implicit reconstruction of the scene with severe blurry images and corresponding event data.
+2) An intensity-change-metric event loss and a photo-metric blur loss are presented to supervise the modeled motion blur in terms of both camera pose change and image formation conjugatively.
+3) Experiments on extended synthetic data and real-captured data validate that our method achieves accurate motion trajectory estimation and high-quality 3D implicit reconstruction of the scene with severe blurry images and corresponding event data.
 2. Related Work
-Table 1. Comparison of Previous Deblurring NeRF Works
-Method RGB Event Motion Blur Model
-Deblur-NeRF - 2D Blur Kernal
-DP-NeRF - 3D Blur Kernal
-Sharp-NeRF - Grid-Based Blur Kernal
-BAD-NeRF - Linear Interpolation Bundle Adjustment
-E2NeRF Poses from Pre-deblurred Images
-EBAD-NeRF Event-driven Bundle Adjustment
 2.1. Neural Radiance Fields
 The neural radiation fields model achieves impressive novel view synthesis results and inspires a lot of subsequent research (Gao et al., 2022).
-For example, (Barron et al., 2021, 2022; Verbin et al., 2022; Xu et al., 2022) improve the quality of learned geometry and synthesized novel views.
-Other works try to improve the training and rendering speed such as FastNeRF (Garbin et al., 2021), MobileNeRF (Chen et al., 2023) and EfficientNeRF (Hu et al., 2022).
-PixelNeRF (Yu et al., 2021) and RegNeRF (Niemeyer et al., 2022) realize reconstructing high-quality NeRF with few input views.
-RawNeRF (Mildenhall et al., 2022) can recover a clean NeRF from noisy raw images.
-HDR-NeRF generates high dynamic novel view images with low dynamic images as input.
-Besides, nerf also requires accurate camera poses from COLMAP as input.
-BARF (Lin et al., 2021), L2G-NeRF, and Nope-NeRF (Bian et al., 2023) explore reconstructing NeRF when camera poses are not provided.
+Some works (Barron et al., 2021, 2022; Verbin et al., 2022; Xu et al., 2022) improve the quality of learned geometry and synthesized novel views.
+Other works try to improve the training and rendering speed (Garbin et al., 2021; Chen et al., 2023; Hu et al., 2022).
+Besides, BARF (Lin et al., 2021), L2G-NeRF, and Nope-NeRF (Bian et al., 2023) explore reconstructing NeRF without camera poses estimated by COLMAP (Schonberger and Frahm, 2016) and input images.
 Blurry Images are often captured in low-light scenes or with a high-speed moving camera on drones or robots.
@@
 However, the blurring kernel is optimized based on 2D-pixel location independently.
-DP-NeRF (Lee et al., 2023) proposes a rigid blurring kernel and generates a 3D deformation field, which is constructed as the 3D rigid motion of the camera for each view.
+DP-NeRF (Lee et al., 2023) proposes a rigid blurring kernel and generates a 3D deformation fields, which is constructed as the 3D rigid motion of the camera for each view.
 Sharp NeRF (Lee et al., 2024) proposed a learnable grid-based kernel to obtain sharp output from neural radiance fields.
@@
 Traditional deblurring algorithms focus on finding the suitable blur kernel of a blurry image to recover a sharp image.
-Hand-crafted features or sparse priors are used to tackle this problem blur (Chan and Wong, 1998; Krishnan et al., 2011; Xu et al., 2013).
-Deep learning deblurring works directly learn an end-to-end mapping from blurry to sharp images (Tao et al., 2018a; Wieschollek et al., 2017; Zhang et al., 2020).
+Hand-crafted features or sparse priors are used to tackle this problem (Chan and Wong, 1998; Krishnan et al., 2011; Xu et al., 2013).
+Deep learning deblurring works directly learn an end-to-end mapping from blurry images to sharp images (Tao et al., 2018a; Wieschollek et al., 2017; Zhang et al., 2020).
 MPRNet (Zamir et al., 2021a) and SRN (Tao et al., 2018b) achieve impressive single-image deblurring results.
-In addition to single-image deblurring, (Kim et al., 2016; Zhang and Carin, 2014; Zhang and Yang, 2015) proposes deblurring schemes for continuous video images, which used the characteristics of video images(such as temporal consistency) to better predict sharp images, but this also brought greater calculation amount.
-He et al.
-(He et al., 2016) weights a combination of input images to avoid explicit calculation of blur kernels and effectively improves processing speed.
-Other works have also tried implementing multiple image deblurring using convolutional neural network and Unet framework (Wieschollek et al., 2016; Su et al., 2017).
-Son et al.(Son et al., 2021) proposes an effective recurrent video deblurring network and achieves state-of-the-art performance.
-However, the information missing from blurry images during exposure time makes the robustness of the image-based method decreases in the face of various motion blur.
+However, the information missing from blurry images during exposure time decreases the robustness of the image-based methods and limits their performance when facing various and severe motion blur in low-light or high-speed scenes.
+Table 1. Comparison of Previous Deblurring NeRF Works
+Method RGB Event Motion Blur Trajectory Modeling
+Deblur-NeRF - 2D Blur kernel
+DP-NeRF - 3D Blur kernel
+Sharp-NeRF - Grid-based Blur kernel
+BAD-NeRF - Linear Interpolation Bundle Adjustment
+E2NeRF Fixed Poses from Pre-deblurred Images
+EBAD-NeRF Event-driven Bundle Adjustment
 2.3. Event Camera
-Event camera is based on a bio=inspired vision senor that measures brightness changing asynchronously (Gallego et al., 2020).
-The high temporal resolution and high dynamic visual data capturing paradigm makes event cameras ideal for flow estimation (Shiba et al., 2022; Hagenaars et al., 2021; Gehrig et al., 2021; Akolkar et al., 2020; Pan et al., 2020), feature detection and tracking(Xu et al., 2020; Zhang et al., 2021, 2022) and images deblurring (Pan et al., 2019; Lin et al., 2020; Shang et al., 2021; Xu et al., 2021).
+The event camera is based on a bio-inspired vision sensor that measures brightness changing asynchronously (Gallego et al., 2020).
+The high temporal resolution and high dynamic visual data capturing paradigm makes it ideal for flow estimation (Shiba et al., 2022; Hagenaars et al., 2021; Gehrig et al., 2021; Akolkar et al., 2020; Pan et al., 2020), feature detection and tracking (Xu et al., 2020; Zhang et al., 2021, 2022) and images deblurring (Pan et al., 2019; Lin et al., 2020; Shang et al., 2021; Xu et al., 2021).
 Pan et al.
 (Pan et al., 2019) propose a simple event-based double integral model (EDI) based on event generation to establish the connection between blurry images and events.
-Shang et al.
-(Shang et al., 2021), Jiang et al.(Jiang et al., 2020) and Sun et al.
-(Sun et al., 2022) using learning-based network to recover sharp images with events.
-Recently, event-based NeRF works have emerged.
+Other works (Shang et al., 2021; Jiang et al., 2020; Sun et al., 2022) use learning-based networks to recover sharp images with events.
+Recently, some event-based NeRF works have emerged.
 Ev-NeRF (Hwang et al., 2023) models the measurement of the event sensor to learn grayscale neural radiance fields derived from the event stream.
@@
 e-NeRF (Low and Lee, 2023) improves Ev-NeRF and EventNeRF for non-uniform camera motion.
-DE-NeRF (Ma et al., 2023) uses both RGB images and events to learn a deformable NeRF.
-E2NeRF first reconstructs a sharp NeRF with blurry images and corresponding event data and achieves state-of-the-art performance.
+SpikeNeRF (Zhu et al., 2024) derives a NeRF-based volumetric scene representation from spike camera data.
+DE-NeRF (Ma et al., 2023) uses RGB images and events to learn a deformable NeRF.
+E2NeRF is the first to reconstruct a sharp NeRF with blurry images and corresponding event data and achieves state-of-the-art performance.
 However, it only explicitly models the blurring processing at the imaging aspect.
-For camera motion during the exposure time, E2NeRF relies on pre-deblurred images with EDI (Pan et al., 2019) algorithm and structure-from-motion method COLMAP for pose estimation.
-Notably, our proposed EBAD-NeRF transforms camera motion trajectories into a learnable model with event-driven bundle adjustment.
-The framework unlocks the potential of event information in explicit modeling of blur processing at the level of both image changes and camera motion.
+For camera motion trajectory during the exposure time, E2NeRF uses EDI (Pan et al., 2019) to pre-deblurred images and COLMAP to estimate the camera poses that can not be optimized during training.
+Ev-DeblurNeRF (Cannici and Scaramuzza, 2024) extends from DP-NeRF (Lee et al., 2023) with a novel EDI loss and an eCRF network to simulate the generation of events more realistically and achieves a remarkable deblurring effect.
+It also uses extra continuous events between spare image views to enhance the NeRF learning, which differs from the setting of E2NeRF and ours.
 Figure 2.
 Overview of our method.
-An image with motion blur is caused by the camera moving during the exposure time.
-We temporal evenly sample $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ and transfer them into SE(3) as learnable parameters.
-Then, the poses are input into the NeRF network to model the blurry image and event stream generation.
-The intensity-change-metric loss can optimize the modeling in terms of camera motion.
-The photo-metric loss can optimize and optimize the modeling in terms of image formation.
-With the joint learning of NeRF parameters $\theta$ and camera poses, a sharp NeRF is learned eventually.
+For each view of the static scene, an image with motion blur is caused by the camera moving during the exposure time.
+We temporal evenly sample $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ on the motion trajectory and transfer them into SE(3) as learnable variables $\{\mathbf{T}_{i}\}_{i=1}^{p}$.
+With the poses, we model the blurry image and event generation in our network and jointly optimize the NeRF parameters $\theta$ and motion trajectory with the ERGB data to learn a sharp NeRF eventually.
 3. Method
-We introduce EBAD-NeRF to simultaneously learn the poses of the camera’s motion process and generate sharp neural radiance fields using both blurred images and the event stream within the corresponding exposure time.
+We introduce EBAD-NeRF to simultaneously learn the poses during the camera motion blur process and sharp neural radiance fields with blurred images and the event stream within the corresponding exposure time.
 Events serve two key purposes within the framework: (1) optimizing camera trajectories within the exposure time, and (2) contributing to the physical formation of motion blur in the image by providing information on light intensity changes.
-Moreover, we employ a photo-metric blur loss to supervise blur formation in the RGB domain.
-Figure 2 shows the overview of our approach.
-Sec.
-3.1 introduces the motion blur formation in the neural radiance fields.
-Sec.
-3.2 illustrate the proposed event-driven bundle adjustment method.
-Sec.
-3.3 shows the loss function and the derivation, and Sec.
-3.4 describes the implementation details of our method.
+Moreover, we employ a photo-metric loss to supervise blur formation in the RGB domain.
+Figure 2 is the overview of our approach.
 3.1. Motion Blur Formation in Static 3D Scene
 In a static 3D scene, camera motion blur is caused by the changes of camera pose $\mathbf{P}$ during the exposure time.
-A digital image $\mathcal{I}$ is obtained by measuring the integration of light intensity $I$ on the image sensor.
+A digital image $\mathcal{I}$ is obtained by measuring the integration of light intensity $I$ with respect to time $t$ on the image sensor:
 (1) $\mathcal{I}=\int_{t_{start}}^{t_{end}}I(t)\mathrm{d}t,\$
-where ${t_{start}}$ and ${t_{end}}$ is the start and end time of exposure.
-The light intensity $I$ is only corresponding to the camera pose, and we can express $\mathbf{P}$ as a function of time $\mathbf{P}=p(t)$.
-Then, a blurry image $\mathcal{B}$ can be expressed as:
+where ${t_{start}}$ and ${t_{end}}$ are the start and end time of the exposure period.
+Since light intensity $I$ is only corresponding to the camera pose in the static scene, we can express $\mathbf{P}$ as a function of time $\mathbf{P}=p(t)$ and a blurry image $\mathcal{B}$ can be expressed as:
 (2) $\mathcal{B}=\int_{t_{start}}^{t_{end}}I(P(t))\mathrm{d}t.\$
-In NeRF (Mildenhall et al., 2021), the rendering results are closely related to camera poses.
-Given a camera pose $\mathbf{P}=P(t)$, for each pixel of the imaging plane, we can obtain a ray $\mathbf{r}$ that emits from the optical center of the camera and passes through the pixel $\mathbf{x}$.
-With stratified sampling, we can divide the part of this ray starting with $l_{near}$ and ending with $l_{far}$ into N equal parts and randomly sample points in each part.
-For each of the $N$ sampled points, we input the 3D coordination $\mathbf{o}$ in the world coordinate system and 2D view the direction $\mathbf{d}$ represented by the ray $\mathbf{r}$ into NeRF MLP $F_{\theta}$ with network parameters $\theta$:
-(3) $(\mathbf{c},\sigma)=F_{\mathbf{\theta}}(\gamma_{o}(\mathbf{o}),\gamma_{d}(\mathbf{d})),$
-the output of which is color $\mathbf{c}$ and density $\sigma$ and $\gamma(\cdot)$ is an encoder for a higher $K+1$ dimension of input of the network:
+In NeRF (Mildenhall et al., 2021), given a camera pose $\mathbf{P}=P(t)$, for each pixel of the imaging plane, we can obtain a ray $\mathbf{r}$ that emits from the optical center of the camera and passes through the pixel $\mathbf{x}=(x,y)$.
+With stratified sampling, we can divide the part of this ray starting with $l_{near}$ and ending with $l_{far}$ into N equal parts and randomly sample one point in each part.
+For each of the $N$ sampled points, we input its 3D coordination $\mathbf{o}$ in the world coordinate system and the 2D view direction $\mathbf{d}$ represented by the ray $\mathbf{r}$ into NeRF MLP $F_{\theta}$ with network parameters $\theta$:
+(3) $(\mathbf{c},\sigma)=F_{\mathbf{\theta}}(\gamma_{o}(\mathbf{o}),\gamma_{d}(\mathbf{d})).$
+The outputs are color $\mathbf{c}$ and density $\sigma$ of the point, and $\gamma(\cdot)$ encodes the input to a higher $K+1$ dimension:
 (4) $\gamma_{K}(x)=\{\sin(2^{k}\pi{x}),\cos(2^{k}\pi{x})\}_{k=0}^{K}.$
@@
 $\text{where}\quad T(i)=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}),$
-we can get the final color value of the pixel $\mathbf{x}$ passed by the ray $\mathbf{r}$ corresponding to the pose $\mathbf{P}$.
-$\delta_{i}=l_{i+1}-l_{i}$ is the distance between adjacent sampled points, and $T_{i}$ is the transparency between $l_{n}$ and sampled point.
-If we assume that the learned NeRF MLP is sharp, we can model the image motion blur by discretizing Eq. (2) with $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ sampled evenly over the exposure time:
-(6) ^$\hat{\mathcal{B}}(\mathbf{x})=\frac{1}{p}\sum_{i=1}^{p}C(P(t),\mathbf{x}),\mathbf{x}\in\mathcal{X},\$
+we can get the final color value $C(\mathbf{r},\mathbf{x})=C(\mathbf{P},\mathbf{x})$ of the pixel $\mathbf{x}$ passed by the ray $\mathbf{r}$ corresponding to the pose $\mathbf{P}$.
+$\delta_{i}=l_{i+1}-l_{i}$ is the distance between adjacent sampled points, and $T_{i}$ is the transparency between $l_{near}$ and the sampled point.
+If we assume that the learned NeRF MLP is sharp, we can model the image motion blur by discretizing Eq.
+(2) with $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}=\{P(t_{i})\}_{i=1}^{p}$ sampled evenly over the exposure time $t_{start}$ to $t_{end}$:
+(6) ^$\hat{\mathcal{B}}(\mathbf{x})=\frac{1}{p}\sum_{i=1}^{p}C(P(t_{i}),\mathbf{x}),\mathbf{x}\in\mathcal{X},\$
 where $\mathcal{X}$ represents the pixels of the image sensor.
-Since we use virtual sharp image color $C(P(t),\mathbf{x})$ to replace the light intensity $I(P(t))$, we need to multiply each item by a weight of time in Eq.
+Since we use virtual sharp image color $C(P(t_{i}),\mathbf{x})$ to replace the light intensity $I(P(t_{i}))$ of Eq.
+(2), we need to multiply each item by a weight of time.
+As the poses are temporally evenly sampled, we can directly use an average as $\frac{1}{p}$ to represent the weight as in Eq.
 (6).
-As the poses are temporally evenly sampled, we can directly use an average $\frac{1}{p}$ to represent the weight.
-At this point, we have established the connection between image motion blur, camera pose, and neural radiation field.
+At this point, we establish a connection between image motion blur, camera pose, and neural radiation fields.
 3.2. Event-driven Bundle Adjustment
-In BAD-NeRF (Wang et al., 2023), only the start pose $\mathbf{p}_{1}$ and end pose $\mathbf{p}_{p}$ of the exposure time is parameterized and learned, and the motion process is represented with a linear model.
-This approach may fail in long exposure or high-speed situations where the camera is likely to have complex motion (non-uniform motion, curved motion) during exposure time.
-In order to better fit the camera motion trajectory, we express the camera poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ as $\{\mathbf{T}_{i}\}_{i=1}^{p}\in\mathbf{SE}(3)$.
-At this time, in addition to recovering the starting and ending poses, the motion trajectory modeling must also find the relationship between adjacent poses.
-We can model this correlation by modeling the generation of events, which measures the light intensity change in the log domain asynchronously:
-(7) $E(t_{1},t_{2})=\left\{\begin{aligned} \lfloor\frac{\log(I(t_{2}))-\log(I(t_{1})))}{\Theta}\rfloor,&I(t_{2})\leq I(t_{1})\\ \lceil\frac{\log(I(t_{2}))-\log(I(t_{1}))}{\Theta}\rceil,&I(t_{2})>I(t_{1})\end{aligned}.\right.$
-Note that we round the modeled events because an event is only triggered when the light intensity change threshold reaches $\Theta$.
-According to previous event simulation work, V2E (Hu et al., 2021), we can convert the color of the images into grayscale and model the event generation.
-Based on this, the estimated event data can be expressed by bringing the predicted color in Eq.
-(5) into Eq.
-(7):
-(8) $\hat{E}(t_{1},t_{2})=\left\{\begin{aligned} \lfloor\frac{\log(\overline{C}(T({t_{2}})))-\log(\overline{C}(T(t_{1})))}{\Theta}\rfloor,&\overline{C}(t_{2})\leq\overline{C}(t_{1})\\ \lceil\frac{\log(\overline{C}(T({t_{2}})))-\log(\overline{C}(T(t_{1})))}{\Theta}\rceil,&\overline{C}(t_{2})>\overline{C}(t_{1})\end{aligned},\right.$
-where $\overline{C}(\cdot)$ means taking the average on the RGB channels to get the intensity value.
-The computation of $\hat{E}(\mathbf{x},t_{1},t_{2})$ is differentiable with respect to $\{\mathbf{T}_{i}\}_{i=1}^{p}$, and we can jointly optimize the sampled camera poses and parameters of $F_{\theta}$ by using actual events as supervision.
-3.3. Loss Functions
-We introduce the intensity-change-metric event loss to optimize the camera motion poses $\{\mathbf{T}_{i}\}_{i=1}^{p}$ during exposure time as well as the NeRF parameter $\theta$ by minimizing the differences between the numbers of predicted and actual events on all pixels $\mathbf{x}\in\mathcal{X}$:
-(9) $\mathcal{L}_{event}=\frac{1}{p}\sum_{k=1}^{p-1}\|(\hat{E}_{k}(t_{k},t_{k+1}))-(E_{k}(t_{k},t_{k+1}))\|_{2}^{2}.$
-In the training phase, we only calculate the event loss between adjacent sampling poses because if the sampling frames are far apart, the event loss will fluctuate greatly, thus affecting the stability of the network.
-We also conduct the photo-metric blur loss between the predicted and the input blurry images:
-(10) $\mathcal{L}_{blur}={\|\hat{B}(\mathbf{x})-B(\mathbf{x})\|}_{2}^{2},$
-which also has a constraint on both the camera poses and neural radiance fields, as proved in BAD-NeRF.
-To strengthen the learning, the final loss is the sum of $\mathcal{L}_{event}$ and $\mathcal{L}_{blur}$:
-(11) $\mathcal{L}_{blur}=\lambda\mathcal{L}_{event}^{c}+\mathcal{L}_{blur}^{c}+\mathcal{L}_{blur}^{f},$
+In long exposure or high-speed situations, the camera is likely to have complex motion (non-uniform motion) during exposure time.
+To better model the motion blur, we transform the sampled $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ into $p$ learnable pose $\{\mathbf{T}_{i}\}_{i=1}^{p}\in\mathbf{SE}(3)$ at time $\{t_{i}\}_{i=1}^{p}$.
+Then, we can use events as a bridge between camera pose changes and neural radiance fields to bundle these poses into actual camera motion trajectory during the NeRF optimization.
+As described in Sec.3.1, we can express the virtual sharp image color during the exposure as $C(\mathbf{T}_{i},\mathbf{x})=C(\mathbf{P}_{i},\mathbf{x})$ and convert it into grayscale as $G(\mathbf{T}_{i},\mathbf{x})$ by averaging the RGB channels.
+Then, according to previous event simulation work (Hu et al., 2021), we transform $G(\mathbf{T}_{i},\mathbf{x})$ into the log function domain as Ł$\L(\mathbf{T}_{i},\mathbf{x})$ and simulate the generation of events caused by camera motion as:
+(7) ^$\hat{\mathcal{E}}(\mathbf{x},t_{i},t_{i+1})=\left\{\begin{aligned} \frac{\L(\mathbf{T}_{i+1},\mathbf{x})-\L(\mathbf{T}_{i},\mathbf{x})}{\Theta},&\L(\mathbf{T}_{i+1},\mathbf{x})\leq\L(\mathbf{T}_{i},\mathbf{x})\\ \frac{\L(\mathbf{T}_{i+1},\mathbf{x})-\L(\mathbf{T}_{i},\mathbf{x})}{\Theta},&\L(\mathbf{T}_{i+1},\mathbf{x})>\L(\mathbf{T}_{i},\mathbf{x})\end{aligned}.\right.$
+Then with the real events $\mathcal{E}(\mathbf{x},t_{i},t_{i+1})$ captured by the event camera as event-driven bundle adjustment, we can optimize adjunctive poses $T_{i}$ and $T_{i+1}$ with the intensity-change-metric event loss by minimizing the differences between the numbers of predicted and actual events on all $m$ pixels $\mathbf{x}\in\mathcal{X}$ among all input views:
+(8) $\mathcal{L}_{event}=\frac{1}{m}\sum_{\mathbf{x}\in\mathcal{X}}\frac{1}{p-1}\sum_{i=1}^{p-1}\|(\hat{\mathcal{E}}(\mathbf{x},t_{i},t_{i+1}))-(\mathcal{E}(\mathbf{x},t_{i},t_{i+1}))\|_{2}^{2}.$
+The computation of ^$\hat{\mathcal{E}}(\mathbf{x},t_{1},t_{2})$ is differentiable with respect to $\{\mathbf{T}_{i}\}_{i=1}^{p}$ and latent sharp NeRF parameters $\theta$, which is optimized by the Jacobians of event loss:
+(9) $\frac{\partial{\mathcal{L}_{event}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\hat{\mathcal{E}}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\hat{\mathcal{E}}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
+(10) $\frac{\partial{\mathcal{L}_{event}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{event}}{\partial\hat{\mathcal{E}}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\hat{\mathcal{E}}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}},$
+Notice that in Eq. (8) we only calculate the event loss between adjacent poses because if the sampled poses are too far apart, the event loss will fluctuate greatly, reducing the training stability.
+3.3. Final Loss
+We conduct the photo-metric blur loss between the predicted and the input blurry images as in NeRF for all $m$ pixels $\mathbf{x}\in\mathcal{X}$ among all input views:
+(11) $\mathcal{L}_{blur}=\frac{1}{m}\sum_{\mathbf{x}\in\mathcal{X}}{\|\hat{\mathcal{B}}(\mathbf{x})-\mathcal{B}(\mathbf{x})\|}_{2}^{2},$
+The blur loss not only has a constraint on camera poses $\{\mathbf{T}_{i}\}_{i=1}^{p}$ to avoid the spreading of poses but also plays a main role in supervising NeRF parameters $\theta$ to learn texture details of the scene that event data cannot provide:
+(12) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\hat{\mathcal{B}}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\hat{\mathcal{B}}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
+(13) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{blur}}{\partial\hat{\mathcal{B}}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\hat{\mathcal{B}}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}}.$
+The final loss is defined as:
+(14) $\mathcal{L}_{blur}=\lambda\mathcal{L}_{event}^{f}+\mathcal{L}_{blur}^{c}+\mathcal{L}_{blur}^{f},$
 where $\lambda$ is a weight parameter of event loss.
-We used the design of the fine and coarse network in NeRF to conduct blur loss $\mathcal{L}^{f}_{blur}$ and $\mathcal{L}^{c}_{blur}$ on both of the networks.
-Event loss $\mathcal{L}^{f}_{event}$ is only calculated for the fine network because it can render enough texture details for event estimation.
-The learnable camera poses $\{\mathbf{T}_{i}\}_{i=1}^{p}$ and NeRF parameter $\theta$ are optimized by the Jacobians of event loss:
-(12) $\frac{\partial{\mathcal{L}_{event}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\mathcal{E}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{E}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
-(13) $\frac{\partial{\mathcal{L}_{event}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{event}}{\partial\mathcal{E}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{E}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}},$
-and the Jacobians of blur loss:
-(14) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\mathcal{B}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{B}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
-(15) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{blur}}{\partial\mathcal{B}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{B}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}},$
-where $m$ is the number of pixels of the image.
-Then, the network can learn an accurate camera motion and sharp 3D implicit representation with gradient propagation.
+We used the design of the fine and coarse network in NeRF and conduct blur loss $\mathcal{L}^{f}_{blur}$ and $\mathcal{L}^{c}_{blur}$ on both of the networks.
+Event loss $\mathcal{L}^{f}_{event}$ is only calculated for the fine network because it can render enough texture details for precise event estimation.
+Then, the network can learn an accurate camera motion and latent sharp 3D implicit representation with gradient propagation.
 3.4. Implementation Details
 We train the EBAD-NeRF on a single NVIDIA RTX 3090 GPU.
-The training time has not increased compared with BAD-NeRF because there is no extra rendering computation for event loss.
-The sample number of the rays $N=64$ for fine and coarse networks.
-The number of sampled poses of each view $p$ is set to 4, and the weight of event loss $\lambda$ is set to 0.005.
+The training time is close to BAD-NeRF.
+The sample number of the rays is $N=64$ for both fine and coarse networks.
+The number of sampled poses of each view $p$ is set to 5, and the weight of event loss $\lambda$ is set to 0.005.
 Sec.4.4 evaluates the influence of the parameters.
@@
 A coarse initial pose for each view is given at the start of training.
-4. EXPERIEMENTS
+4. Experiments
 4.1. Datasets
 4.1.1. Synthetic Data:
-We extend five Blender scenarios of Deblur-NeRF (Cozyroom, Factory, Pool, Tanabata, and Wine).
-(1) We increase the camera shake amplitude of the Factory, Pool, Tanabata, and Wine scenes to increase the blur degree of the training images and maintain the camera shaking in Deblur-NeRF of the Cozyroom scene to evaluate the effect of different shaking scales.
-(2) During the camera motion, we added the speed changes often occurring when the hand-held camera shakes.
-(3) To generate corresponding event data, we input the virtual sharp frames of the camera shake process rendered by Blender into the event simulation tool V2E (Hu et al., 2021) to generate the corresponding event stream.
+We use the five Blender scenarios of Deblur-NeRF (Cozyroom, Factory, Pool, Tanabata, and Wine) to generate our synthetic data.
+We increase the camera shake amplitude to generate a severe blur and add the camera motion speed change during the exposure to simulate a non-uniform camera motion blur for the Factory, Pool, Tanabata, and Wine scenes.
+Besides, we input the virtual sharp frames during the camera shake process rendered by Blender into the V2E (Hu et al., 2021) to generate the corresponding event stream as in the synthetic datasets of E2NeRF.
 4.1.2. Real Data:
-To test the effectiveness of our method on real captured data, we captured two sets of real data using the DAVIS-346 (Taverni et al., 2018) event camera, which can capture spatial-temporal aligned event data and RGB data at the same time.
-Both sets of data are captured in low-light scenes, so a longer exposure time (100ms) is needed for RGB data to get a bright enough image.
-We use a handheld camera to capture training data with blur and a tripod to mount the camera to capture sharp testing data.
-Each scene contains 16 views of blurry images and corresponding events for training and 4 novel view sharp images for training.
-4.2. Comparing Method
-4.2.1. Image Deblurring Method:
+To test the effectiveness of our method on real data, we capture two sets of real data (Bar and Classroom) with DAVIS-346 (Taverni et al., 2018) event camera, which can capture spatial-temporal aligned event data and RGB data at the same time.
+Both sets of data are captured in low-light scenes, so RGB data requires a longer exposure time (100ms) to get a bright enough image.
+We use a handheld camera to capture blurry images and events for training and a tripod to capture sharp images for testing.
+Each scene contains 16 views of blurry images and corresponding events for training and 4 novel view sharp images for testing.
+4.2. Comparing Methods and Metrics
+4.2.1. Image Deblurring Methods
 For comparison, we selected two classic learning-based image deblurring methods, MPR (Zamir et al., 2021a) and SRN (Tao et al., 2018b).
-In order to make a more fair comparison in terms of data modalities, we selected the D2Net method, which also uses event and image modal data fusion to achieve image deblurring.
-We input the images deblurred by these three methods into the original NeRF for the novel view generation task to learn a sharp 3D representation.
-4.2.2. Deblurring NeRF Method:
-We selected image-based deblurring NeRF method Deblur-NeRF (Ma et al., 2022) and BAD-NeRF (Wang et al., 2023) and ERGB-dual-modal-based deblurring NeRF method E2NeRF (Qi et al., 2023) for comparison.
-For Deblur-NeRF, we use the default parameters for the spare deformation kernel.
-For BAD-NeRF, we use the number of interpolated virtual cameras as 7 to get sufficiently satisfactory results, as described in the paper.
-For E2NeRF, we follow the paper and use $b=4$ to train the model.
+In order to make a more fair comparison in terms of data modalities, we use D2Net (Shang et al., 2021), which also uses events and blurry image data to achieve image deblurring.
+We input the images deblurred by these three methods into the original NeRF for the novel view generation task and named them as MPR+NeRF, SRN+NeRF, and D2Net+NeRF.
+4.2.2. Deblurring NeRF Methods:
+We selected image-based deblurring NeRF methods Deblur-NeRF (Ma et al., 2022) and BAD-NeRF (Wang et al., 2023) and ERGB deblurring NeRF method E2NeRF (Qi et al., 2023).
 All NeRF-based methods are trained with 5000 rays of batch size and 64 sampling points for coarse and fine networks.
+Other parameters are the default values of the methods.
 We perform 100,000 iterations on synthetic data.
-On real data, due to the reduction in image resolution and the number of views, 50,000 iterations are enough for all NeRF-based methods to converge to good enough results.
+On real data, due to the reduction in image resolution and the number of input views, 50,000 iterations are enough for all NeRF-based methods to converge to the final results.
 4.2.3. Metrics:
-We use PSNR, SSIM, and LPIPS (Zhang et al., 2018) to evaluate the NeRF reconstruction results and use absolute trajectory error (ATE) to evaluate the fitting accuracy of the camera motion trajectory (Grupp, 2017).
+We use PSNR, SSIM, and LPIPS (Zhang et al., 2018) to evaluate the reconstruction results and use absolute trajectory error (ATE) to evaluate the fitting accuracy of the camera motion trajectory (Grupp, 2017).
 Table 2. Quantitative Ablation Study on the Tanabata Scene.
-PSNR↑ SSIM↑ LPIPS↓ ATE
+PSNR↑ SSIM↑ LPIPS↓ ATE↓
 BAD-NeRF 20.65 .7567 .3349 0.0514 $\pm$ 0.0204
@@
 Figure 4.
-Ablation study of trajectory fitting accuracy on Tanabata scene.
+Qualitative ablation study of trajectory fitting accuracy on Tanabata scene.
 EBAD-NeRF-linear and EBAD-NeRF-noe are defined in Sec.
 4.3.
-The figure shows that the event-driven bundle adjustment significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than other methods.
+EBAD-NeRF significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than all other methods.
+4.3. Event-driven Bundle Adjustment Analysis
+We analyze the proposed event-driven bundle adjustment at two levels: 3D reconstruction quality and motion trajectory fitting accuracy.
+In Table 2 and Figures 3, 4, EBAD-NeRF-noe training the proposed EBAD-NeRF without event data as supervision and EBAD-NeRF-linear represents the results of adding event data constraints to BAD-NeRF with linear interpolation directly.
+Results in Table 2 are the average of the deblurring view and novel view.
+4.3.1. Defect of Linear Interpolation Bundle Adjustment:
+The quantitative results in Table 2 demonstrate that without event enhancement, the bundle adjustment constrained by linear interpolation (BAD-NeRF) is even worse than the unconstrained bundle adjustment (EBAD-NeRF-noe) when facing complex camera motion.
+With event enhancement, the same conclusion can be obtained by comparing EBAD-NeRF-linear and EBAD-NeRF-full.
+4.3.2. Effect of Event in Bundle Adjustment:
+Comparing BAD-NeRF and EBAD-NeRF-linear, even with the negative impact of linear pose interpolation, EBAD-NeRF-linear still has improvements in trajectory fitting and reconstruction results with event enhancement as shown in Table 2.
+Comparing EBAD-NeRF-noe and EBAD-NeRF, introducing events significantly improves the accuracy of trajectory estimation and reconstruction results.
+To sum up, during the NeRF training, event data can not only provide image-level optimization information but also pose-level practical constraints.
+Our proposed EBAD-NeRF effectively utilizes the ERGB-dual-modal data to achieve this goal.
+The qualitative comparison in Figure 3 and Figure 4 also testify to this, where EBAD-NeRF generates the sharpest rendering results and closest trajectory estimation to the ground truth.
+4.4. Parameters Analysis
+We evaluate the rendering results of the Cozyroom scene to analyze the influence of $p$ and $\lambda$ of our method in Figure 5.
+4.4.1. The influence of $p$:
+Intuitively, the more virtual sharp frames, the more accurate the simulation of camera motion blur will be, and the corresponding network learning results will be better.
+The results in Figure 5 are consistent with this.
+When the number of virtual frames gradually increases from 1, the network’s performance also significantly improves.
+However, after exceeding 5, the number of virtual frames has almost no impact on the reconstruction results, and more virtual sharp frames will require more calculations.
+Therefore, $p=5$ is set in our experiments.
+4.4.2. The influence of $\lambda$:
+As shown in Figure 5, when we take $\lambda$ as 0.005, EBAD-NeRF achieves the best results on the three metrics.
+Figure 5.
+Evaluation on the influence of the number of the sampled poses $p$ and the weight parameter $\lambda$.
+The results are averages of reconstructed images on both deblurring and novel views of the Cozyroom scene.
+The red, green, and blue lines represent PSNR, SSIM, and LPIPS, respectively.
+Table 3. Absolute Trajectory Error Results of Blender Scenes.
+ATE↓
+COLMAP-Blur (NeRF (Mildenhall et al., 2021)) 0.0476 $\pm$ 0.0171
+COLMAP-EDI (E2NeRF (Qi et al., 2023)) 0.0419 $\pm$ 0.0159
+BAD-NeRF (Linear Interpolation(Wang et al., 2023)) 0.0462 $\pm$ 0.0142
+BAD-NeRF (Cubic Interpolation(Wang et al., 2023)) 0.0542 $\pm$ 0.0207
+Event-driven Bundle Adjustment (Ours) 0.0383 $\pm$ 0.0143
 Table 4. Quantitative Results on Deblurring Views of Blender Scenes. The Best Results are Shown in Bold.
@@
 BAD-NeRF 28.80 .9278 .1872 20.39 .6696 .4052 29.46 .8867 .2416 20.35 .7475 .3363 22.10 .7540 .3831 24.22 .7971 .3107
+BAD-NeRF-Cubic 28.89 .9314 .1664 25.63 .8526 .2822 30.26 .9049 .2078 21.31 .7677 .3706 23.22 .7923 .3384 25.86 .8498 .2731
 E2NeRF 30.17 .9459 .1057 27.90 .9046 .2638 29.29 .8841 .2420 24.02 .8625 .2906 25.63 .8688 .2919 27.40 .8932 .2388
-EBAD-NeRF 30.53 .9475 .1120 28.10 .9085 0.159 31.50 .9193 .1607 24.91 .8783 .2071 26.66 .8732 .2230 28.30 .9031 .1749
-4.3. Event-driven Bundle Adjustment Analysis
-We analyze the proposed event-driven bundle adjustment at two levels: 3D reconstruction quality and motion trajectory fitting accuracy.
-As shown in Table 2, Figure 3 and Figure 4, EBAD-NeRF-noe training proposed EBAD-NeRF without event data as supervision.
-EBAD-NeRF-linear represents the results of adding event data constraints to BAD-NeRF directly.
-4.3.1. Defect of Linear Interpolation Bundle Adjustment:
-The quantitative results in Table 2 demonstrate that without event enhancement, the bundle adjustment constrained by linear interpolation (BAD-NeRF) is even worse than the unconstrained bundle adjustment (EBAD-NeRF-noe) for fitting irregular camera motions.
-The same conclusion will be obtained by comparing EBAD-NeRF-linear and EBAD-NeRF-full, both with event enhancement.
-4.3.2. Effect of Linear Event in Bundle Adjustment:
-Comparing BAD-NeRF and EBAD-NeRF-linear, even with the negative impact of linear interpolation bundle adjustment in Table 2, EBAD-NeRF-linear still has improvements in trajectory fitting and reconstruction results.
-Comparing EBAD-NeRF-linear and EBAD-NeRF under the constraints of event information, event-driven bundle adjustment achieved the best trajectory fitting and reconstruction results.
-To sum up, during the NeRF training, event data can not only provide image-level optimization information but also pose-level practical constraints.
-Our proposed EBAD-NeRF effectively utilizes the ERGB-dual-modal data to achieve this goal.
-The qualitative comparison in Figure 3 and Figure 4 also testify to this.
-Figure 5.
-The influence of the number of learnable poses $p$ and the weight parameter $\lambda$ on the Cozyroom scene.
-The results are average for reconstructed images from both deblurring and novel views.
-4.4. Analysis on Parameters
-We evaluate the rendering results on the Cozyroom scene to analyze the superparameters $p$ and $\lambda$ of the proposed network.
-The red, green, and blue lines in Figure 5 represent PSNR, SSIM, and LPIPS, respectively.
-4.4.1. The influence of $b$:
-Intuitively, the more virtual sharp frames, the more accurate the simulation of camera motion blur, and the corresponding network learning results will be better.
-The results in Figure 5 are consistent with this.
-When the number of virtual frames gradually increases from 1, the performance of the network also becomes significantly better.
-After exceeding 5, the number of virtual frames has almost no impact on the reconstruction results.
-Therefore, $p=5$ is set in our experiments.
-4.4.2. The influence of $\lambda$:
-As shown in Figure 5, $\lambda$ has the best effect when set to 0.005, and if the weight of event loss is further increased, the network performance will gradually decrease.
+EBAD-NeRF 30.53 .9475 .1120 28.10 .9085 .1717 31.50 .9193 .1607 24.91 .8783 .2071 26.66 .8732 .2230 28.34 .9054 .1749
 Table 5. Quantitative Results on Novel Views of Blender Scenes. The Best Results are Shown in Bold
@@
 BAD-NeRF 28.52 .9250 .1918 21.11 .7169 .4077 29.70 .8919 .2376 22.37 .8102 .3267 22.79 .7791 .3812 24.90 .8246 .3090
+BAD-NeRF-Cubic 29.14 .9337 .1677 23.25 .7758 .3179 30.68 .9134 .2047 20.37 .7432 .4260 23.82 .8185 .3358 25.45 .8369 .2904
 E2NeRF 30.07 .9469 .1062 27.78 .9090 .2654 29.34 .8905 .2374 24.25 .8730 .2882 25.70 .8772 .2910 27.43 .8993 .2376
-EBAD-NeRF 30.60 .9476 .1138 28.25 .9146 .1583 31.62 .9244 .1559 25.45 .8898 .2056 26.72 .8889 .1961 27.80 .9131 .1784
+EBAD-NeRF 30.60 .9476 .1138 28.25 .9146 .1583 31.62 .9244 .1559 25.45 .8898 .2056 26.72 .8889 .1961 28.53 .9131 .1659
 Table 6. Quantitative Results on Real Scenes. The Results are the Average of the Scenes, and the Best Results are Shown in Bold.
-NeRF NeRF+D2Net NeRF+MPR NeRF+SRN Deblur-NeRF BAD-NeRF E2NeRF EBAD-NeRF
-PSNR↑ 24.87 26.31 24.35 25.91 25.59 28.10 27.78 29.60
-SSIM↑ .8562 .8829 .8539 .8706 .8744 .8839 .9062 .9175
-LPIPS↓ .5005 .4092 .4963 .4198 .3885 .3632 .2403 .2156
+NeRF NeRF+D2Net NeRF+MPR NeRF+SRN Deblur-NeRF BAD-NeRF BAD-NeRF-Cubic E2NeRF EBAD-NeRF
+PSNR↑ 24.87 26.31 24.35 25.91 25.59 28.10 27.03 27.78 29.60
+SSIM↑ .8562 .8829 .8539 .8706 .8744 .8839 .8744 .9062 .9175
+LPIPS↓ .5005 .4092 .4963 .4198 .3885 .3632 .2978 .2403 .2156
 4.5. Quantitative Results
+4.5.1. Quantitative Results of Pose Estimation on Blender Scenes:
+We use the Evo tool (Grupp, 2017) to calculate the absolute trajectory error, considering both rotation and translation.
+The results in Table 3 are the average of five blender scenes, which shows the accuracy of camera motion trajectory estimation by different methods.
+COLMAP-Blur means directly estimating the camera poses with blurry images and COLMAP for each view.
+COLMAP-EDI represents the pose estimation method used in E2NeRF, which uses the EDI (Pan et al., 2019) algorithm to first deblur the image with events and then input the sharp image sequence into COLMAP to estimate camera poses during the blur process.
+BAD-NeRF learns the start and end poses of camera motion.
+It uses linear interpolation to simulate the intermediate poses and extends it with a cubic interpolation.
+The results of these two methods are also shown in the table.
+Our proposed event-driven bundle adjustment method introduces event into joint learning of camera pose changes and sharp NeRF, achieving the best results in restoring the camera pose, which also promotes EBAD-NeRF to further improve the reconstruction quantification results as in Sec.
+4.5.2 and Sec.
+4.5.3.
 4.5.2. Quantitative Results of Reconstruction on Blender Scenes:
-In Tables 3 and 4, we evaluate the reconstruction results of learned neural radiance fields in five blender scenarios from deblurring and novel views, respectively.
-SRN and SRN+NeRF achieve the best results in single-image-based deblurring methods.
-With the help of event data, D2Net is also slightly better than MPR.
-For Deblur-NeRF and BAD-NeRF, although they have been further improved compared to NeRF+D2Net, NeRF+MPR, and NeRF+SRN, they still cannot surpass event-enhanced E2NeRF and our proposed EBAD-NeRF.
-Note that there is no apparent difference between the E2NeRF and EBAD-NeRF on the Cozyroom scene where linear camera shake produces motion blur, and E2NeRF is even better than EBAD-NeRF on LPIPS.
+We evaluate the reconstruction results of five blender scenarios on both deblurring views (Table 4) and novel views (Table 5).
+SRN and SRN+NeRF achieve the best results among single-image-deblurring-based methods and even surpass the Deblur-NeRF.
+With the help of event data, D2Net and D2Net+NeRF are slightly better than MPR and MPR+NeRF.
+BAD-NeRF with cubic interpolation is further improved over BAD-NeRF with linear interpolation and SRN+NeRF, though it is still worse than event-enhanced methods E2NeRF and our proposed EBAD-NeRF on both novel view and deblurring view.
+Note that there is no apparent difference between the E2NeRF and EBAD-NeRF on the Cozyroom scene where blur is only caused by linear camera motion, and E2NeRF is even better than EBAD-NeRF on LPIPS.
 However, in the other four scenes with more complex camera shaking as described in Sec.
-4.1, the results of EBAD-NeRF in both deblurring views and novel views are significantly better than E2NeRF, which indicates that lacking the optimization of poses limits the role of events in 3D reconstruction in E2NeRF, especially when facing complex camera motion.
+4.1.1, the results of EBAD-NeRF in both deblurring views and novel views are significantly better than E2NeRF, which indicates that with fixed pre-estimated poses limits the effect of events in 3D implicit learning in E2NeRF, especially when facing complex camera motion.
 Accordingly, EBAD-NeRF uses event-driven bundle adjustment to jointly optimize the motion poses of the camera, further releasing the potential of event data in reconstructing a sharp NeRF with blurry images.
-Table 6. Absolute Trajectory Error Results of Blender Scenes.
-ATE
-COLMAP-Blur (NeRF (Mildenhall et al., 2021)) 0.0476 $\pm$ 0.0171
-COLMAP-EDI (E2NeRF (Qi et al., 2023)) 0.0419 $\pm$ 0.0159
-Linear Bundle Adjustment (BAD-NeRF (Wang et al., 2023)) 0.0462 $\pm$ 0.0142
-Event-driven Bundle Adjustment (Ours) 0.0383 $\pm$ 0.0143
-4.5.2. Quantitative Results of Pose Estimation on Blender Scenes:
-In Table 6, we use absolute trajectory error to evaluate the accuracy of camera motion trajectory estimation by different methods.
-The results in the table are the average of five blender scenes.
-COLMAP-Blur means directly inputting the blurry images into COLMAP to estimate the camera poses of each view.
-This is a typical pose estimation method in NeRF-related work.
-COLMAP-EDI represents the method used in E2NeRF.
-It uses the EDI (Pan et al., 2019) algorithm to first use events to deblur the image and then input the sharp image sequence into COLMAP to estimate camera poses of the blur process.
-Linear bundle adjustment represents the linear interpolation method used by BAD-NeRF to jointly learn the pose changes of each view.
-Our proposed event-driven bundle adjustment method introduces event into joint learning of camera pose changes and sharp NeRF, achieving the best results in restoring the camera pose, which also promotes EBAD-NeRF to further improve the reconstruction quantification results as in Tables 3 and 4.
-4.5.3. Quantitative Reconstruction Results on Blender Scenes:
-To verify the effectiveness of our method on real data, we conducted experiments with two real scenes captured by DAVIS-346.
-As shown in Table 5, the comparison results with other methods are highly consistent with the results on blender scenes in Tables 3 and 4, which verifies the reliability of our method in practical applications.
-Figure 6. Qualitative results of deblurring views on blender Pool scene.
-Figure 7. Qualitative results of novel views on blender Wine and Factory scenes (first and second row), real scene (third row).
+4.5.3. Quantitative Results of Reconstruction on Real Scenes:
+To verify the effectiveness of our method on real data, we conducted experiments on real data mentioned in 4.1.2.
+As shown in Table 6, the comparison results with other methods are highly consistent with the conclusions obtained in Tables 4 and 5, which verifies the reliability of our method in practical applications.
 4.6. Qualitative Results
 4.6.1. Qualitative Results of Reconstruction on Blender Scenes
-As shown in Figure 6 and the first and second row in Figure 7, we compared the results of the synthesis deblurring view on the pool scene and novel view on the Wine and Factory scenes, respectively.
-For the deblurring view results, the results of EBAD-NeRF are sharper than other state-of-the-art single-image deblurring methods, event-RGB-based deblurring methods, and deblurring NeRF methods.
-For the green plants and wooden boxes in the scene, some artifacts appear in the results of E2NeRF, and our method effectively recovers the texture details.
+We evaluate the reconstruction results of the deblurring view on the Pool scene (Figure 6) and the novel view on the Wine and Factory scenes (first two rows in Figure 7).
+Since image-deblurring methods cannot synthesize novel view images, we do not show their results in Figure 7).
+The results of EBAD-NeRF are sharper than all other methods on both deblurring and novel views.
+For the green plants and wooden boxes in the Pool scene, some artifacts appear in the results of E2NeRF, and our method effectively recovers the texture details.
 In the Wine and Factory scenes, though E2NeRF realizes noticeable results for the characters and letters, our method is closer to ground truth in color, clarity, and line thickness.
+Figure 6. Qualitative deblurring views rendering results of blender Pool scene.
+Figure 7. Qualitative novel views rendering results of blender scenes (first and second row) and real scenes (third row).
 Figure 8.
-Qualitative evaluation of trajectory fitting accuracy on Tanabata scene.
-The figure shows that the event-driven bundle adjustment significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than other methods.
-Note that though there is event enhancement on blurry image formation in E2NeRF, the motion trajectory is distorted and leads to the inaccurate NeRF.
+Qualitative evaluation of trajectory estimation on Tanabata scene.
+EBAD-NeRF significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than all other methods.
+Although E2NeRF uses pre-deblurred images with events, the estimated motion trajectory is still distorted, leading to reconstruction degradation even with event enhancement during training.
+4.6.2. Qualitative Results of Reconstruction on Real Scenes
+As in the third row of Figure 7, our method better recovers the white letters on the blackboard compared to E2NeRF.
+BAD-NeRF is limited in reconstructing smooth areas (white areas on the podium).
+The results of other image-deblurring-NeRF methods and Deblur-NeRF are all limited by the severe blurry input, which is consist with the quantitative results in Table 6.
 4.6.3. Qualitative Results of Pose Estimation on Blender Scenes
 In Figure 8, we compare the estimated camera trajectory of different methods from the three dimensions of roll, pitch, and yaw.
-The purple curve representing EBAD-NeRF is the closest to the dashed line representing ground truth, which is consistent with the quantitative data in Table 6.
-4.6.3. Qualitative Results on Real Scenes
-As in the third row of Figure 7, our method better recovers the white letters on the blackboard, compared to the pose-fixed event-RGB-based method E2NeRF, while BAD-NeRF is limited in reconstructing smooth areas (white areas on the podium).
-The results of other image-deblurring-NeRF methods and Deblur-NeRF are all limited by the severe blurry input, which is the same as the quantitative results in Table 5.
-Experiments on synthetic and real data show that our method can learn better poses of the camera motion blur process with event-driven bundle adjustment compared to the previous deblurring NeRF method.
-On this basis, the event data is also superimposed on the image blur process supervision, eventually achieving better 3D reconstruction and novel view synthesis effects.
+The purple curve representing EBAD-NeRF is the closest to the ground truth dashed line, which is consistent with the results in Table 3.
+Experiments on synthetic and real data both show that our method can learn better poses of the camera motion blur process with event-driven bundle adjustment compared to the previous deblurring NeRF methods.
+Upon this basis, the event data is also superimposed on the image blur process supervision, eventually achieving better 3D reconstruction and novel view synthesis effects.
 5. Conclusion
-In conclusion, this paper presents a novel approach, leveraging event-driven bundle adjustment, to address the challenge of modeling camera motion blur within neural radiance fields (NeRFs).
-Our framework draws inspiration from the emerging field of event-based vision, which offers high temporal resolution and asynchronous events, ideal for capturing dynamic scenes with motion blur.
+In conclusion, this paper presents a novel approach, leveraging event-driven bundle adjustment, to address the challenge of modeling camera motion blur within neural radiance fields (NeRF).
+We draw inspiration from the emerging field of event-based vision, which offers high temporal resolution events, ideal for capturing dynamic information in motion blur.
 By integrating intensity-change-metric event loss and photo-metric blur loss, our framework enables the simultaneous optimization of blur modeling alongside NeRF reconstruction.
-Experiments on synthetic and real-world datasets demonstrate the efficacy of our approach in accurately estimating camera poses and producing sharp NeRF reconstructions.
+Experiments on synthetic and real-captured datasets demonstrate the efficacy of our approach in accurately estimating camera poses and producing sharp NeRF reconstruction results.
+Appendix A Supplementary Qualitative Comparison on Synthetic Scenes
+As shown in Figure 9, in the Cozyroom scene there is no apparent qualitative difference in the results of the BAD-NeRF (Wang et al., 2023), E2NeRF (Qi et al., 2023) and EBAD-NeRF because the camera motion is even and slight.
+However, in the other four scenes, the results of our EBAD-NeRF have been significantly improved compared to BAD-NeRF and E2NeRF, proving that when facing complicated camera motion, our proposed event-driven bundle adjustment can effectively provide accurate estimated poses for the NeRF network and with event-enhanced imaging simulation our method can recover a sharp NeRF from the severely blurry images and corresponding events.
+In the Factory scene, EBAD-NeRF’s reconstruction of scars on the wall is closest to ground truth.
+In comparison, the results of E2NeRF are too thick, and the results of BAD-NeRF are still blurry.
+In the Pool scene, E2NeRF misses the lotus leaf in the pool and generates blurrier results in the Tanabata scene than EBAD-NeRF.
+Artifacts appear on the white flag in the results of E2NeRF in the Wine scene.
+Since there is no supervision of event data for pose estimation and image motion blur modeling, the results of BAD-NeRF are not sharp enough compared to ERGB-based methods E2NeRF and EBAD-NeRF.
+The pre-deblurring-based method MPR-NeRF (Zamir et al., 2021a), SRN-NeRF (Tao et al., 2018b), and D2Net-NeRF (Shang et al., 2021) is not good enough due to the limitations of the deblurring method itself.
+Overall, our EBAD-NeRF achieved the best results in all synthetic scenes.
+Appendix B Supplementary Qualitative Comparison on Real-World Scenes
+Figure 10 shows the two scenes of the real data.
+In the Bar scene, the synthesis novel view results of EBAD-NeRF are closest to ground truth, especially for the light spot in the upper left corner and the characters on the billboard.
+Other methods have obvious reconstruction errors of the light points and blur results of the characters.
+In the Classroom scene, our EBAD-NeRF obtains sharper letters and fewer artifacts in the synthesis results than other methods.
+The experiments on real-world scenes demonstrate that our EBAD-NeRF is more effective than the ERGB-based deblurring NeRF method E2NeRF, which without pose optimizing during the training.
+Also, our method has significantly improved over the image-based deblurring NeRF methods and NeRF with pre-deblurring images.
+Appendix C Supplementary Video
+We show the video rendering results for the ablation study, synthetic scenes, and real-world scenes on our project website: https://icvteam.github.io/EBAD-NeRF.html.
+The video rendering results for the ablation study prove the effectiveness of our proposed event-driven bundle adjustment.
+For the synthetic scene, our method achieves the sharpest video rendering results.
+The results of E2NeRF are a little misaligned in space because the poses estimated by its pre-deblurring framework are slightly different from the initial poses given for the other methods.
+For the real-world scene, the results of NeRF and NeRF with pre-deblurring images are severely blurry and contain many cloudy materials.
+In comparison, our EBAD-NeRF has sharper rendering results with less cloudy materials than Deblur-NeRF, BAD-NeRF, and E2NeRF.
+Overall, our EBAD-NeRF achieves the best video rendering results in both real-world and synthetic scenes.
+Figure 9. Supplementary Reconstruction Qualitative Comparison on Synthetic Scenes.
+Figure 10. Supplementary Reconstruction Qualitative Comparison on Real Scenes.